Write a digit in the blank space of each of the following numbers so that the number formed is divisible by 11: 
 8 __ 9484

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let ‘a’ be placed here
Sum of its digits at odd places = 4 + 4 + a
=8 +a
Sum of its digits at even places = 8 + 9 + 8
= 25
Difference = 25–(8 + a)
= 17–a
The difference should be 0 or a multiple of 11, then the number is divisible by 11
If 17–a = 0
a = 17 (which is not possible)
Now, take a multiple of 11.
Let’s take 1117–a = 11
a = 17–11
a=6
Therefore the required digit is 6

Thanks